# Decomposed Prompting: Unveiling Multilingual Linguistic Structure Knowledge in English-Centric Large Language Models

### v1

@@ Abstract
Despite the predominance of English in their training data, English-centric Large Language Models (LLMs) like GPT-3 and LLaMA display a remarkable ability to perform multilingual tasks, raising questions about the depth and nature of their cross-lingual capabilities. This paper introduces the decomposed prompting approach to probe the linguistic structure understanding of these LLMs in sequence labeling tasks. Diverging from the single text-to-text prompt, our method generates for each token of the input sentence an individual prompt which asks for its linguistic label. We assess our method on the Universal Dependencies part-of-speech tagging dataset for 38 languages, utilizing both English-centric and multilingual LLMs. Our findings show that decomposed prompting surpasses the iterative prompting baseline in efficacy and efficiency under zero- and few-shot settings. Further analysis reveals the influence of evaluation methods and the use of instructions in prompts. Our multilingual investigation shows that English-centric language models perform better on average than multilingual models. Our study offers insights into the multilingual transferability of English-centric LLMs, contributing to the understanding of their multilingual linguistic knowledge.
Decomposed Prompting: Unveiling Multilingual Linguistic Structure Knowledge in English-Centric Large Language Models
Ercong Nie1,2, Shuzhou Yuan3, Bolei Ma2,4, Helmut Schmid1, Michael Färber3, Frauke Kreuter2,4,5 and Hinrich Schütze1,2 1Center for Information and Language Processing (CIS), LMU Munich, 2Munich Center for Machine Learning (MCML), 3Karlsruhe Institute of Technology, 4Department of Statistics, LMU Munich, 5University of Maryland, College Park nie@cis.lmu.de, shuzhou.yuan@kit.edu, bolei.ma@lmu.de,
@@ 1 Introduction
Current Large Language Models (LLMs), such as GPT-3, GPT-4, PaLM, and LLaMA (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023a), have demonstrated remarkable capabilities in in-context learning, also known as prompting, across a broad spectrum of language understanding and generation tasks (Zhao et al., 2023; Zhang et al., 2023; Ziyu et al., 2023). These models are predominantly trained on massive amounts of English text data, with some limited exposure to other languages. For instance, LLaMA2’s pretraining corpus comprises over 89% English content (Touvron et al., 2023b). Yet, these English-centric LLMs 111In this paper, we regard a model pretrained primarily on English text as English-centric. still exhibit effective performance in multilingual evaluations (Lai et al., 2023). In a multilingual prompting scenario designed for zero-shot transfer with LLMs, the model executes tasks by directly generating outputs based on a task description and/or a few examples provided in a pivot language (typically English), along with input in a different target language (Ahuja et al., 2023). However, the extent and nature of their cross-lingual capabilities remain underexplored (Ye et al., 2023). This raises a critical question: Does the multilinguality of these models stem from a deep, generalizable multilingual linguistic understanding, or merely from the superficial alignment of lexical patterns across languages?
Given the demonstrated proficiency of English-centric LLMs in multilingual tasks that demand profound language understanding (Deng et al., 2023; Wang et al., 2023b), we hypothesize that these models harbor substantial multilingual knowledge. This knowledge, particularly relating to linguistic structure, is commonly conceptualized through sequence tagging tasks (Jurafsky, 2000). However, the current prompting strategies designed for sequence labeling in LLMs are not well suited to test. For instance, behavioral probing methods (Belinkov et al., 2020), aimed at measuring knowledge stored in language models, struggle to adapt to tasks predicting more complex structures. Additionally, text-to-text prompting methods (Asai et al., 2023), which rely on a predefined output template, face challenges in maintaining control over the output format. In response to these challenges, a decent iterative prompting strategy for structured prediction has been introduced, addressing the aforementioned limitations (Blevins et al., 2023). Despite its advantages, this method presents its own challenges, such as longer processing times due to its iterative inference strategy.
@@
To overcome the challenges identified in probing the multilingual knowledge of linguistic structure in LLMs, we introduce the decomposed prompting strategy. This strategy is inspired by the ToPro method (Ma et al., 2024), a novel prompt-based fine-tuning approach for sequence labeling tasks. We adopt this idea to the in-context learning paradigm, aiming to probe English-centric LLMs for their understanding of token-level linguistic structure framed as sequence labeling tasks. As shown in Figure 1, instead of employing a single text-to-text prompt for labeling an entire sequence in one step, our method decomposes this process into multiple discrete prompts. More precisely, we first split the input sentence into tokens. Subsequently, we generate an individual prompt for each token which inquires about its linguistic label.
@@
We evaluate our approach on the Universal Dependency (UD) part-of-speech (POS) tagging dataset (Nivre et al., 2020) covering 38 languages with 3 English-centric LLMs and 2 multilingual LLMs. Our approach outperforms the iterative prompting baseline in both zero- and few-shot settings in terms of accuracy and efficiency. We investigate the nuanced impact of evaluation methods and the usage of task instructions within prompts on the performance of decomposed prompting, followed by an empirical comparative study of decomposed and iterative prompting. Moreover, our analysis of multilingual efficacy of English-centric LLMs yields valuable insights into the transferability of linguistic knowledge via multilingual prompting.
@@ 2 Background and Related Work
English-Centric LLMs
English-centric LLMs are primarily pretrained on large English text data. Despite this focus, these models are also exposed to a limited amount of multilingual data. LLaMA (Touvron et al., 2023a), for example, is pretrained on an extensive scale of corpora comprising over 1.4 trillion tokens, of which less than 4.5% constitute multilingual data from 20 different languages. LLaMA 2 (Touvron et al., 2023b) expands this linguistic diversity, featuring 27 languages each representing more than 0.005% of the pertaining data. Mistral 7B (Jiang et al., 2023), a cutting-edge English-centric LLM, achieves superior performance and efficiency through the adoption of advanced attention techniques such as Sliding Window Attention (SWA) (Child et al., 2019), facilitating faster inference. To enhance the robustness of multilingual processing, the Byte-level Byte-Pair-Encoding (BBPE) algorithm (Sennrich et al., 2016; Wang et al., 2020) is commonly used for tokenization in LLMs. This approach is able to decompose UTF-8 characters, which are outside the scope of the model vocabulary, into their constituent bytes. Consequently, BBPE tokenization equips LLMs with the versatility to handle scripts from any language, theoretically, even those not encountered during training. These two factors—limited exposure to non-English data and byte-level encoding capability—jointly contribute to the robust multilingual abilities observed in English-centric LLMs.
Multilingual LLMs and Instruction Tuning
LLMs are typically instruction-tuned models, including multilingual LLMs, as exemplified by models such as BLOOMZ (Muennighoff et al., 2023), which is derived from BLOOM (Workshop et al., 2022), and mTk (Wang et al., 2022) which is based on mT5 (Xue et al., 2021). Instruction tuning is a widely utilized method to improve the task understanding and interactivity of LLMs (Zhang et al., 2023). Recent studies augmented the multilingual capabilities of LLMs by applying the multilingual instruction tuning (Kew et al., 2023; Chen et al., 2023; Shaham et al., 2024). Multilingual LLMs have also been tailored for specific language clusters, such as the SeaLLMs for Southeast Asian Languages (Nguyen et al., 2023). In our work, we utilize instruction-tuned English-centric and multilingual LLMs for our experiments.
@@
Prompting methods have seldom been applied to sequence labeling tasks. However, some studies have used prompt-based fine-tuning for this purpose. Cui et al. (2021) have implemented template-based prompting techniques with the BART model for Named Entity Recognition (NER) tasks. Ma et al. (2022) introduced a template-free prompting strategy for few-shot NER, termed entity-oriented LM fine-tuning. Similar to our approach, Ma et al. (2024) utilized a decomposition-based prompting method to fine-tune multilingual encoder models for cross-lingual sequence labeling. Contrary to these studies, which all operate within a fine-tuning framework, our research is distinctively focused on facilitating multilingual in-context learning in LLMs without the necessity for parameter updates. Prompting LLMs for sequence labeling tasks remains a challenge (Ahuja et al., 2023). While text-to-text prompting is widely adopted across various benchmarking tasks for LLMs (Lai et al., 2023), their application to sequence labeling is hindered by the intricate requirements of output templates (Asai et al., 2023), which may not accurately capture the full capabilities of LLMs. Iterative structured promoting, specifically designed for sequence labeling, has been introduced recently (Blevins et al., 2023). In this approach, the model decodes in step a label for the word at position of the sequence. This predicted label, along with the next word, is then input back into the model to predict the next label. The dependency of each token’s prediction on the preceding one substantially slows down the inference process. In contrast, our proposed decomposed prompting method offers improvements in both efficacy and efficiency.
@@ 3 Decomposed Prompting for LLMs
In this study, we introduce a novel approach for conducting sequence labeling with LLMs through in-context learning, termed decomposed prompting.
3.1 Intuition
3.2 Problem Formulation
@@
In decomposed prompting, we design a prompt template function which generates a specific prompt for each token. takes the input sequence and an individual token as arguments and returns a prompt for predicting the label of the token. The true label can be optionally included as an argument to ; if included, utilizes to provide a demonstration. An example of such a template function is illustrated as follows.
@@
4 Experimental Setup
4.1 Datasets and Languages
@@
In our study, we focus on evaluating the multilingual linguistic structure knowledge of English-centric models through multilingual part-of-speech tagging tasks, employing our proposed decomposed prompting method. We utilize a subset of the Universal Dependency treebanks (UDPOS) (Nivre et al., 2020) for this purpose. The UDPOS dataset adopts a universal POS tag set consisting of 17 tags (Appendix A.1). Our chosen subset, derived from the XTREME multilingual benchmark (Hu et al., 2020), comprises 38 languages from diverse language families distributions (Appendix A.2). If the test set of a language contains more than 200 sentences, we randomly sample 200 instances for the evaluation due to computational constraints.
4.2 Models
We select a diverse list of LLMs, including three English-centric LLMs and two multilingual LLMs in order to investigate the differences in multilingual understanding across LLMs with varying degrees of multilinguality and base capabilities. All LLMs in this experiment are instruction-tuned versions accessible through the HuggingFace framework (Wolf et al., 2020).
English-centric LLMs
LLaMA2 represents an advanced iteration of the LLaMA foundation models developed by Meta AI (Touvron et al., 2023a, b), trained on publicly available corpora predominantly in English. We consider LLaMA2 models with 7B and 13B parameters in our experiments. Mistral 7B (Jiang et al., 2023) enhances the LLaMA models in terms of both performance and inference efficiency, achieved through meticulous engineering in language model design and training. For our experiments, we utilize the instruction-tuned version of Mistral 7B, which has been fine-tuned on the OpenHermes 2.5 dataset 222https://huggingface.co/datasets/teknium/OpenHermes-2.5.
Multilingual LLMs
BLOOMZ (Muennighoff et al., 2023) is a multi-task fine-tuned variant of the BLOOM model (Workshop et al., 2022), which is trained on 46 languages. We employ its 7B version in our experiment. mTk-Instruct (Wang et al., 2022) is a multilingual encoder-decoder model, fine-tuned on instruction-following datasets. It is built upon the mT5 model (Xue et al., 2021), which is pretrained on corpora of over 100 languages. It comprises approximately 13 billion parameters.
4.3 Baselines and Settings
Iterative Prompting (Iter)
Decomposed Prompting (Decom)
To evaluate our proposed approach, we employ the prompt template outlined in §3.2 to decompose the entire sequence into a set of individual prompts for prediction. In our experiments, we use the 17 POS tags themselves as the label words, i.e., we expect the model to directly predict a tag from the tagset shown in Appendix A.1 by selecting the tag with the highest logit.
Zero- and Few-Shot Prompting
We devised two experimental scenarios for multilingual prompting—zero-shot and few-shot—to evaluate the performance of both approaches under different conditions. In the zero-shot setting, only an English task instruction is provided alongside the input in the target language. The text in Figure 6 of Appendix A.1, which outlines the tag set information, serves as the instruction in our experiments. In few-shot prompting, we supplement the prompt with a few English demonstrations, structured according to the prompt template of each method. For Decom, we randomly select an example for each tag type from the English training set to create a demonstration. For a fair comparison, the same number of demonstrations are used for the Iter baseline. We refer to Appendix B for the details of prompts used in the experiments.
Evaluation Methods
We contrast two evaluation methodologies for prompting in our experiments. The probability-based method leverages the model’s output logits to retrieve the probability distribution over the tag set, subsequently identifying the label with the highest probability. In case the label word is tokenized into subtokens, we use the first subtoken to serve as the label word, following previous work (Zhao et al., 2021; Wang et al., 2023a). The generation-based method directly compares the content generated by the LLM with the gold label.
We use the weighted average F1 scores for different tags as our evaluation metric. All experiments were conducted on a server equipped with 4 A100-SXM4-80GB GPUs.
| Zero-shot | Few-shot | Avg. | |||
| en | mult. | en | mult. | ||
| LLaMA2-7B | |||||
| Iter (prob.) | 33.1 | 27.2 | 68.0 | 48.6 | 44.2 |
| Decom (prob.) | 58.2 | 43.2 | 74.7 | 50.5 | 56.7 |
| Decom (gen.) | 53.8 | 40.4 | 62.1 | 45.8 | 50.5 |
| LLaMA2-13B | |||||
| Iter (prob.) | 47.6 | 37.4 | 68.0 | 52.6 | 51.4 |
| Decom (prob.) | 67.3 | 54.7 | 77.3 | 54.5 | 63.5 |
| Decom (gen.) | 59.2 | 48.7 | 65.3 | 48.3 | 55.4 |
| Mistral-7B | |||||
| Iter (prob.) | 65.2 | 54.3 | 80.2 | 58.9 | 64.7 |
| Decom (prob.) | 63.6 | 61.8 | 85.0 | 64.4 | 68.7 |
| Decom (gen.) | 45.3 | 48.7 | 81.4 | 63.0 | 59.6 |
| BLOOMZ | LLaMA2 | Mistral | Avg. | |
@@ 4.3 Baselines and Settings
5 Results and Analysis
We evaluate the performance of iterative and decomposed prompting for English and multilingual POS tag labeling tasks under zero- and few-shot settings. Our goal is (1) to validate the benefits of decomposed prompting in comparison to the baseline method (§5), and (2) to explore the extent to which decomposed prompting captures multilingual linguistic structure knowledge from the English-centric LLMs (§6).
5.1 Main Findings
The overall results for English-centric LLMs, as detailed in Table 1, demonstrate that our proposed decomposed prompting obviously outperforms the iterative prompting baseline across both zero- and few-shot settings, in both English and multilingual evaluations. This trend holds true for all three English-centric models tested, with the sole exception in the zero-shot setting for the English evaluation with the Mistral-7B model, where Decom slightly lags behind Iter (63.6 vs. 65.2). In addition to superior performance, decomposed prompting offers enhanced efficiency during inference, especially with few-shot prompting. As demonstrated in Table 2, our proposed method achieves, on average, a 2.4-fold increase in speed compared to the baseline in the zero-shot prompting setting and a 6.7-fold increase in the few-shot setting. The efficiency advantage is less obvious with Mistral, owing to Mistral’s implementation of a modified attention mechanism designed to enhance inference efficiency.
5.2 Ablation Study
We performed an ablation study to investigate two factors: the evaluation method used and the type of instructional prompts. Figure 3 presents the outcomes of the ablation study across various model architectures and experimental settings.
| Case 1 |
| Input: Die Lage mitten im Niederdorf , wo Abends am meisten los ist , ist wirklich sensationel . |
| Gloss: the location middle in Niederdorf , where evenings at most happening is , is really sensational . |
| True: DET NOUN ADV ADP PROPN PUNCT ADV ADV ADP PRON ADV VERB PUNCT AUX ADV ADJ PUNCT |
| Iter: DET NOUN ADV ADP PROPN PUNCT PRON PROPN ADP ADP NOUN AUX PUNCT VERB ADP ADJ PUNCT |
| Decom: DET NOUN ADV ADP PROPN PUNCT ADV PROPN ADP ADP ADV VERB PUNCT PRON ADP ADJ PUNCT |
| Case 2 |
| Input: Damit bestätigten die beiden Konzerne Gerüchte , die seit gestern weltweit die Börsianer in Unruhe versetzten . |
| Gloss: thus confirmed the both corporations rumors , which since yesterday worldwide the stock-participants in unrest put . |
| True: ADV VERB DET ADJ NOUN NOUN PUNCT PRON ADP ADV ADJ DET NOUN ADP NOUN VERB PUNCT |
| Iter: ADP VERB PRON ADJ PROPN NOUN PUNCT PRON ADP ADP ADP PRON PROPN ADP NOUN VERB PUNCT |
| Decom: ADV VERB PRON PRON NOUN NOUN PUNCT PRON ADP ADV ADV PRON NOUN ADP NOUN VERB PUNCT |
Probability-Based vs. Generation-Based
We observe in Figure 3(a) that the probability-based approach consistently outperforms the generation-based method for both LLaMA2 versions and the Mistral model. This trend is evident in both the English and multilingual tasks, under zero-shot and few-shot conditions. Generally, the performance in few-shot conditions is better than in zero-shot conditions. Notably, in the Mistral-7B model, the gap between the probability-based and generation-based methods narrows in the few-shot condition.
Effect of Instruction
Figure 3(b) shows that the inclusion of an instruction in prompts has a variable impact across different models and evaluation methods. In probability-based evaluation, the presence of an instruction leads to a noticeable decrease in F1 scores for all models in both English and multilingual tasks. In generation-based evaluation, we also observe some performance decrease in most cases. This suggests that the LLMs better understand the linguistic structure task from the demonstrations than from a task description in natural language.
5.3 Case Study
The case study presented in Table 3 offers a revealing comparative analysis of the iterative and decomposed prompting methods, using selected German examples processed with the Mistral model.
In Case 1, we observe error propagation in the iterative prompting method. The iterative approach incorrectly tags the word “los” as a noun, which subsequently appears to affect the tagging of “ist” as an auxiliary verb rather than the correct tag (VERB), as the tag of the copula verb “ist” depends on the constituent linked to it. This illustrates a fundamental weakness of the iterative method: a single misprediction can adversely influence subsequent predictions and lead to a series of errors.
Case 2 reveals a limitation inherent to decomposed prompting, specifically when a word appears multiple times with varying syntactic functions. Take the word “die” in the given case as an example. This word alternates between a definite article and a relative pronoun. The current design of decomposed prompting lacks the sophistication to discern the distinct grammatical roles that a recurring word can play, consequently assigning the tag PRON for all instances of “die”. Notably, iterative prompting does not resolve this issue in the given example either. This observation underscores the challenge of developing a prompting mechanism capable of context-sensitive discrimination, an area where both decomposed and iterative prompting methods are yet to evolve.
6 Multilinguality Investigation
6.1 Multilingual Performance

### v2

# Decomposed Prompting: Probing Multilingual Linguistic Structure Knowledge in Large Language Models

@@ Abstract
Probing the multilingual knowledge of linguistic structure in LLMs, often characterized as sequence labeling, faces challenges with maintaining output templates in current text-to-text prompting strategies. To solve this, we introduce a decomposed prompting approach for sequence labeling tasks. Diverging from the single text-to-text prompt, our prompt method generates for each token of the input sentence an individual prompt which asks for its linguistic label. We test our method on the Universal Dependencies part-of-speech tagging dataset for 38 languages, using both English-centric and multilingual LLMs. Our findings show that decomposed prompting surpasses the iterative prompting baseline in efficacy and efficiency under zero- and few-shot settings. Moreover, our analysis of multilingual performance of English-centric LLMs yields insights into the transferability of linguistic knowledge via multilingual prompting.
Decomposed Prompting: Probing Multilingual Linguistic Structure Knowledge in Large Language Models
Ercong Nie1,2, Shuzhou Yuan3, Bolei Ma2,4, Helmut Schmid1, Michael Färber3, Frauke Kreuter2,4,5 and Hinrich Schütze1,2 1Center for Information and Language Processing (CIS), LMU Munich, 2Munich Center for Machine Learning (MCML), 3ScaDS.AI and TU Dresden, 4Department of Statistics, LMU Munich, 5University of Maryland, College Park nie@cis.lmu.de, shuzhou.yuan@tu-dresden.de, bolei.ma@lmu.de,
@@ 1 Introduction
Current Large Language Models (LLMs), such as GPT-3, GPT-4, PaLM, and LLaMA (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023a), have demonstrated remarkable capabilities in in-context learning across a broad spectrum of language understanding and generation tasks (Zhao et al., 2023; Zhang et al., 2023; Ziyu et al., 2023). These models are predominantly trained on massive amounts of English text data, with some limited exposure to other languages. For instance, LLaMA2’s pretraining corpus comprises over 89% English content, with the rest in other languages or code (Touvron et al., 2023b). Yet, these English-centric LLMs 111In this paper, we regard a model pretrained primarily on English text as English-centric. exhibit effective performance in complex multilingual language understanding tasks (Deng et al., 2023; Wang et al., 2023). In multilingual evaluation with prompting, a model performs tasks by directly generating outputs based on a task description and/or a few examples provided in a pivot language (typically English), along with input in a different target language (Ahuja et al., 2023). Despite the remarkable multilingual performance of LLMs, the extent and nature of their cross-lingual capabilities remain underexplored (Ye et al., 2023).
@@
We hypothesize that these models harbor substantial multilingual knowledge. This knowledge, particularly relating to linguistic structure, is commonly conceptualized through sequence tagging tasks (Jurafsky, 2000). However, the current prompting strategies designed for sequence labeling in LLMs are not well suited to test. For instance, behavioral probing methods (Belinkov et al., 2020), aimed at measuring knowledge stored in language models, struggle to adapt to tasks predicting more complex structures. To overcome the challenges in probing the multilingual knowledge of linguistic structure in LLMs characterized as sequence labeling, drawing inspiration from the token-level prompt-based fine-tuning method by Ma et al. (2024), we introduce the decomposed prompting strategy, aiming to probe English-centric LLMs for their understanding of linguistic structure framed as sequence labeling tasks. As shown in Figure 1, instead of employing a single text-to-text prompt for labeling an entire sequence in one step, our method decomposes this process into multiple discrete prompts. More precisely, we first split the input sentence into tokens. Subsequently, we generate an individual prompt for each token which inquires about its linguistic label.
@@
We evaluate our approach on the Universal Dependency (UD) part-of-speech (POS) tagging dataset (Nivre et al., 2020) covering 38 languages with 3 English-centric LLMs and 2 multilingual LLMs. Our approach outperforms the iterative prompting baseline in both zero- and few-shot settings in terms of accuracy and efficiency. Furthermore, our investigation into the multilingual performance of English-centric LLMs offers valuable insights into their capabilities of transferring linguistic knowledge through multilingual prompting.
@@ 2 Background and Related Work
Multilinguality of English-Centric LLMs
English-centric LLMs are primarily pretrained on large English text data, with a limited exposion to multilingual data. LLaMA (Touvron et al., 2023a), for example, is pretrained on an extensive scale of corpora comprising over 1.4 trillion tokens, of which less than 4.5% constitute multilingual data from 20 different languages. LLaMA 2 (Touvron et al., 2023b) expands this linguistic diversity, featuring 27 languages each representing more than 0.005% of the pertaining data. Mistral 7B (Jiang et al., 2023) achieves superior performance and efficiency through the adoption of advanced attention techniques such as Sliding Window Attention (SWA) (Child et al., 2019), facilitating faster inference. To enhance the robustness of multilingual processing, the Byte-level Byte-Pair-Encoding (BBPE) algorithm (Sennrich et al., 2016; Wang et al., 2020) is commonly used for tokenization in LLMs. This approach is able to decompose UTF-8 characters, which are outside the scope of the model vocabulary, into their constituent bytes. Thus, BBPE tokenization equips LLMs with the versatility to handle scripts from any language, theoretically, even those not encountered during training. In summary, limited exposure to non-English data and byte-level encoding capability, these two factors discussed above, jointly contribute to the robust multilingual abilities observed in English-centric LLMs.
@@
Prompting LLMs for sequence labeling tasks remains a challenge (Ahuja et al., 2023). While text-to-text prompting is widely adopted across various benchmarking tasks for LLMs (Lai et al., 2023), their application to sequence labeling is hindered by the challenges in maintaining the output templates (Asai et al., 2023). In response, a decent iterative prompting strategy for structured prediction has been introduced (Blevins et al., 2023) (Figure 1). In this approach, the model decodes in step a label for the word at position of the sequence. This predicted label, along with the next word, is then input back into the model to predict the next label. However, the dependency of each token’s prediction on the preceding one substantially slows down the inference process. In contrast, our proposed decomposed prompting method offers improvements in both efficacy and efficiency. Our method is similar to Ma et al. (2024) in that both methods decompose an input sentence into a series of prompts; however, their method is used for fine-tuning, while our method is in an in-context learning paradigm without training.
@@ 3 Decomposed Prompting for LLMs
In decomposed prompting, we design a prompt template function which generates a specific prompt for each token. takes the input sequence and an individual token as arguments and returns a prompt for predicting the label of the token. The true label can be optionally included as an argument to ; if included, utilizes to provide a demonstration.
@@
4 Experiment and Results
4.1 Experimental Setup
Dataset and Language
@@ 4.1 Experimental Setup
We use a subset of the Universal Dependency treebanks (UDPOS) (Nivre et al., 2020) to probe the multilingual linguistic knowledge of LLMs. The UDPOS dataset adopts a universal POS tag set consisting of 17 tags (Appendix A.1.1). Our chosen subset, derived from the XTREME multilingual benchmark (Hu et al., 2020), comprises 38 languages from diverse language families distributions (Appendix A.1.2). We randomly sample 200 instances of each language for the evaluation.
Model and Setup
We experiment on three English-centric LLMs: LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023b), and Mistral-7B (Jiang et al., 2023), as well as two multilingual LLMs: BLOOMZ-7B (Muennighoff et al., 2023) and mTk-Instruct (Wang et al., 2022). All LLMs in our experiment are instruction-tuned versions accessible through the HuggingFace framework (Wolf et al., 2020). We use the weighted average F1 scores for different tags as our evaluation metric. All experiments were conducted on a server with 4 A100-SXM4-80GB GPUs. More details of experimental settings are described in Appendix A.2.
4.2 Overall Results
| Model | Method | Zero-shot | Few-shot | Avg. | ||
| en | mult. | en | mult. | |||
| LLaMA2-7B | Iter | 33.1 | 27.2 | 68.0 | 48.6 | 44.2 |
| Decom | 58.2 | 43.2 | 74.7 | 50.5 | 56.7 | |
| LLaMA2-13B | Iter | 47.6 | 37.4 | 68.0 | 52.6 | 51.4 |
| Decom | 67.3 | 54.7 | 77.3 | 54.5 | 63.5 | |
| Mistral-7B | Iter | 65.2 | 54.3 | 80.2 | 58.9 | 64.7 |
| Decom | 63.6 | 61.8 | 85.0 | 64.4 | 68.7 | |
| BLOOMZ-7B | Decom | 20.6 | 17.6 | 44.1 | 36.2 | 29.6 |
| mTK-Instruct | Decom | 47.6 | 43.1 | 57.3 | 44.7 | 48.2 |
We evaluate the performance of iterative prompting, the baseline method, and decomposed prompting, our proposed method, for English and multilingual POS tag labeling tasks under zero- and few-shot settings. The few-shot examples and the prompts employed in our experiment are presented in Appendix B.2 for reference. Our preliminary experiment to explore the influence of the number of few-shot samples () reveals a mild impact on performance once increases to around 10. More details are provided in Appendix C.
Superiority in Efficacy
The overall results for English-centric LLMs, as detailed in Table 1, demonstrate that our proposed decomposed prompting obviously outperforms the iterative prompting baseline across both zero- and few-shot settings, in both English and multilingual evaluations. This trend holds true for all three English-centric models tested, with the sole exception in the zero-shot setting for the English evaluation with the Mistral-7B model, where Decom slightly lags behind Iter (63.6 vs. 65.2). In addition, English-centric LLMs outperform multilingual LLMs by a considerable margin. The complete experimental results are displayed in Appendix D.
| BLOOMZ-7B | LLaMA2-7B | Mistral-7B | Avg. | |
@@ 4.2 Overall Results
Superiority in Efficiency
In addition to superior performance, decomposed prompting offers enhanced efficiency during inference, especially in few-shot prompting. As demonstrated in Table 2, our proposed method achieves, on average, a 2.4-fold increase in speed compared to the baseline in the zero-shot prompting setting and a 6.7-fold increase in the few-shot setting. The efficiency advantage is less obvious with Mistral, owing to Mistral’s implementation of a modified attention mechanism designed to enhance inference efficiency.
5 Multilingual Analysis